OpenStudy (anonymous):

How do I add/subtract radians when there is only one in the denominator? For example: -18/pi - pi/2

OpenStudy (anonymous):

for an exact number you would just leave it in that form but why would you want to do that? you sure you got the right numbers?

OpenStudy (anonymous):

I am just trying to simplify an equation for a normal to a curve. I am at this point in the problem: y = 6x/pi -18/pi - pi/2

OpenStudy (anonymous):

Get a common denominator of 2π \[-\frac{ 18 }{ \pi }-\frac{ \pi }{ 2 }=-\frac{ 36 }{ 2 \pi }-\frac{ \pi^2 }{ 2 \pi }=\frac{ 36-\pi^2 }{ 2\pi }\]

OpenStudy (anonymous):

but like @090909090909 suggested, I'd leave it like you have it.

OpenStudy (anonymous):

Thank you so much!!!